OpenStudy (anonymous):

Can anyone tell me if I am figuring this problem out right please? Consider the following equation: 6x-9y= -15 A) Write the above equation in the form y=mx+b.. Enter the values of m and b in the appropriate spaces below as integers or reduced fractions (in the form A/B.) Answer: y= ? x + b=? B) Use your answer in part (A) to find the ordered pair that lies on this line when x=45 Answer: (45,? )

OpenStudy (anonymous):

Ok so far I got: 6x-9y= -15 -6x -6x -9y= -6x-15/ -9 I got y=2/3 x + 5/3 Is this correct? & How do I find the coordinated pair? (45, ?)

OpenStudy (anonymous):

put the value of x in the equation

OpenStudy (anonymous):

y=2/3 (45) + 5/3 ? @sritama

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So how do I solve that to get y int? @sritama

OpenStudy (anonymous):